OpenStudy (anonymous):

Variable y varies directly with x2, and y = 96 when x = 4. What is the value of y when x = 2?

OpenStudy (triciaal):

varies directly so y = Kx^2 where k is a constant y/x^2 = k for 2 sets of values x1, y1 and x2, y2 k will be the same y1/x1^2 = y2/x2^2 let (x1, y1) = (4, 96) and (x2, y2) = (2, y) substitute and solve for y

OpenStudy (triciaal):